# ProLex: A Benchmark for Language Proficiency-oriented Lexical Substitution

### v1

@@ Abstract
Lexical Substitution discovers $appropriate$ substitutes for a given target word in a context sentence. However, the task fails to consider substitutes that are of equal or higher proficiency than the target, an aspect that could be beneficial for language learners looking to improve their writing. To bridge this gap, we propose a new task — language proficiency-oriented lexical substitution. We also introduce ProLex, a novel benchmark designed to assess systems’ ability to generate not only appropriate substitutes but also substitutes that demonstrate better language proficiency. Besides the benchmark, we propose models that can automatically perform the new task. We show that our best model, a Llama2-13B model fine-tuned with task-specific synthetic data, outperforms ChatGPT by an average of 3.2% in F-score and achieves comparable results with GPT-4 on ProLex [^1] We will make ProLex publicly available upon publication of the paper..
@@ 1 Introduction
The beginner L2 learners can leverage existing lexical substitution systems Zhou et al. (2019); Lacerra et al. (2021); Yang et al. (2022); Wada et al. (2022); Qiang et al. (2023) to enhance their vocabulary breadth. These systems are designed to identify contextually suitable substitutes for a target word, thereby assisting learners in discovering appropriate substitutions. However, merely knowing these contextually appropriate substitutes is insufficient for L2 learners. Prior benchmarks McCarthy and Navigli (2007); Kremer et al. (2014); Horn et al. (2014); Lee et al. (2021) for lexical substitution focus solely on the contextual appropriateness of generated substitutes. To enhance the vocabulary diversity and writing proficiency of English learners, the substitutions should also reflect an equal or higher level of language proficiency compared to the target word.
In this work, we present ProLex, a novel lexical substitution benchmark that evaluates system performances on language proficiency-oriented lexical substitution, a new task that proposes substitutes that are not only contextually suitable but also demonstrate advanced-level proficiency. To construct ProLex, we begin by selecting target words and sentences according to their frequency in TOEFL-11 Blanchard et al. (2013), a comprehensive corpus of essays written by non-native English learners. This approach ensures a data distribution that is more representative of L2 English learners. Subsequently, to effectively harness human expertise in annotating advanced proficiency-oriented lexical substitutions, we adopt the methodology recommended by Lee et al. (2021) — directly having humans judge the appropriateness of candidate substitutes instead of generating them spontaneously. To facilitate this, we employ GPT-4 to generate an initial pool of substitute candidates. Then, the human annotators assess these candidates based on their contextual appropriateness using our language-learning-oriented annotation scheme. The proposed scheme takes into account the preservation of semantic meaning, common English collocations, lexical diversity, and grammatical correctness. We specifically recruited two annotators who are pursuing their PhDs in Linguistics, specializing in English language teaching and education. Following human annotation, we use the Common European Framework of Reference (CEFR) of Europe Education Committee Modern Languages Division Council for Cultural Co-operation (2001) to remove substitutions that have lower proficiency levels compared to the original target word.
@@ 2 Related Work
The large-scale non-native English writing corpus, TOEFL-11 Blanchard et al. (2013), contains writing samples from non-native TOEFL (the Test of English as a Foreign Language) test takers from 2006 to 2007. In particular, it contains $12,100$ essays, each of which is categorized into one of the three score levels: high, medium, and low. Focusing on enhancing the writing skills of English learning beginners, our methodology involves selecting context sentences and target words primarily from essays with medium and low scores. This approach is based on the assumption that beginner L2 English learners are likely to possess a more limited vocabulary range Fan (2000).
@@ 3.1 Data Creation
The TOEFL-11 dataset comprises of $12,100$ essays, each scored as high, medium, or low. Specifically, there are $4,202$ high-scored essays, $6,568$ with medium scores, and $1,330$ categorized as low. Given the limited vocabulary range of low and medium-proficiency L2 English learners, we intend to select the words that are frequently used in their essays as the target words in our benchmark. For each essay of interest, we select mainly the content words (i.e. nouns, verbs, adjectives, and adverbs) that are used at least three times in the essay. Then, for each such target word, we randomly choose one sentence from all sentences containing that target word. The general statistics of all target words and context sentences are presented in Appendix C Table 5. In sum, we selected $2,531$ unique word-context $(w,s)$ pairs from low and medium-level essays from TOEFL-11 corpus.
@@
We observed that the chosen sentences frequently contain fundamental grammar mistakes, such as spelling errors, which can obscure the intended meaning of the context. Example grammar errors are demonstrated in appendix D. To correct these errors, we follow Rothe et al. (2021) and fine-tune a GPT-2 model Radford et al. (2019) on CLang-8 dataset [^2] https://github.com/ google-research-datasets/clang8. The experiment details are also shown in appendix D. The resulting model is deployed to correct the basic grammar errors in all of our $2,531$ selected context sentences $s$. We then sample $1,125$ $(w,s)$ pairs for human annotations [^3] We used $125$ pairs for calculating inter-annotator agreement and $1,000$ for the annotation. We will include more pairs to expand the benchmark in future work..
@@ 4 Dataset annotation process and statistics
Considering the necessity for substantial semantic understanding and knowledge of English collocations in the annotation process, we recruited two annotators currently pursuing PhDs in Linguistics, with specializations in English language teaching and education. On the $125$ $(w,s)$ pairs, namely $625$ $(w,s,w^{\prime})$ triples, the two annotators reached an inter-annotator agreement of $\kappa=0.6$, indicating substantial agreement [^5] We perform Cohen’s Kappa Cohen (1960) to calculate the inter-annotator agreement. We exclude the labels with 0 to consider only the contexts that are conceivable to both annotators (i.e. labeled as 1 or -1).. Then, they annotated the complete test set separately, with each annotating $2,500$ $(w,s,w^{\prime})$ triples sampled from section 3.1. In total, among all $5,000$ candidate substitutes $w^{\prime}$, $39\%$ is labeled as 1 (accept), while $51\%$ and $10\%$ are annotated as -1 (reject) and 0 (uncertain), respectively. This entails that GPT-4 generated candidates may not always be valid in ProLex, indicating future research opportunities in using LLMs for generating semantically correct, collocationally appropriate English lexical substitutions.
@@
We further process the annotations to compose the final quadruplets $(w,s,w^{a},w^{a}_{p})$ in ProLex, where $w^{a}$ denotes a list of acceptable substitutes for $w$, and $w^{a}_{p}$ denotes the list of acceptable substitutes after filtering based on CEFR levels. We exclude the target words that have zero acceptable substitutes. Table 1 provides an overview of the statistics for the resulting data, grouped by low and medium-level proficiency. In total, there are $680$ word-sentence $(w,s)$ pairs in ProLex with at least one acceptable substitute  [^6] The Part-Of-Speech distribution for the target word $w$ in ProLex is shown in Table 6.. On average, for each $(w,s)$ pair, there are $2.9$ acceptable substitutes $w^{a}$ and $2.3$ proficiency-oriented substitutes $w^{a}_{p}$. Furthermore, we provide the distribution of CEFR levels of the target words in both low and medium sentences, and the result is shown in Figure 2. As evident, a significant proportion of the chosen target words are from A1 and A2 levels, comprising 75% in low-level sentences and 55% in medium-level sentences, respectively. Moreover, in Figure 3, we also present the CEFR level distributions for acceptable substitutes $w^{a}$ and proficiency-oriented substitutes $w^{a}_{p}$. The distribution resonates closely with our research objective of ProLex, which aims to propose acceptable substitutions of high-proficiency level words for commonly used low-proficiency level words in L2 English sentences.
@@ 5.2 Evaluation Metrics
Models BERT-LS Para-LS Llama-2 Vicuna-1.5 ChatGPT GPT-4 7B 13B 7B 13B $D_{S}$ $D_{L}$ $D_{S}$ $D_{L}$ $D_{S}$ $D_{L}$ $D_{S}$ $D_{L}$ $zero$ $32$ $zero$ $32$ Soft $F^{10}$ 25.2 30.2 44.0 48.7 44.5 51.1 44.0 47.8 45.5 52.0 50.9 51.5 50.5 54.7 $F^{10}_{p}$ 20.3 25.3 39.1 43.3 38.6 46.3 38.8 43.0 40.0 46.8 45.2 45.9 44.6 48.8 Hard $F^{10}$ 20.0 20.2 31.4 46.7 27.3 49.1 31.9 45.5 32.1 49.5 47.8 48.3 48.8 52.7 $F^{10}_{p}$ 15.8 17.2 28.1 41.2 24.0 44.2 27.9 40.7 28.5 44.4 42.5 43.0 43.3 46.8
@@
Models Vicuna-1.5 Llama-2 7B 13B 7B 13B Soft $F^{10}$ 49.9 52.1 49.7 54.2 $F^{10}_{p}$ 45.0 46.8 44.5 48.4 Hard $F^{10}$ 48.1 50.2 48.0 51.7 $F^{10}_{p}$ 43.1 44.9 42.7 45.9
@@ 5.3 Baselines
As supervised instruction tuning has proliferated in building powerful instruction following models Ouyang et al. (2022); Bai et al. (2022); OpenAI (2023); Chiang et al. (2023), we experiment with two instruction-tuned large-scale language models, namely Vicuna [^8] We used Vicuna v1.5 provided by Zheng et al. (2023). Chiang et al. (2023) and Llama-2 Touvron et al. (2023). Note that we mainly focus on the smaller variants: Vicuna 7B/13B and Llama-2 7B/13B. Since there is no existing training data for our language proficiency-oriented lexical substitution task, we adopt two approaches to synthesize the data: 1) directly generate data with GPT-4 through prompting, and 2) modify the existing lexical substitution benchmark, Swords. The prompts we employed to synthesize the data with GPT-4 are shown in section E.2.1. As for the second approach, we considered only the acceptable substitutes (i.e. score greater than 50%) from the Swords dataset and filtered the substitutes based on their CEFR levels, by only keeping the ones that have equal or higher proficiency levels compared to the target word. As a result, we denote the training data synthesized with GPT-4 as $D_{L}$, the data modified from Swords as $D_{S}$, and the aggregate of these two datasets as $D_{LS}$.
@@
Leveraging the potent zero-shot capabilities of Large Language Models (LLMs), we evaluate their performance on ProLex by prompting two widely recognized LLMs, GPT-4 and ChatGPT  [^9] Note that we used the version gpt-3.5-turbo-1106, to generate potential substitution candidates. Moreover, we also perform in-context learning with examples selected from $D_{L}$ [^10] We randomly select $32$ examples from $D_{L}$ for in-context learning with LLMs. See appendix E.1 for more details.
@@ 6 Conclusion and Future Work
We propose a new task language proficiency-oriented lexical substitution to improve the vocabulary diversity and writing proficiency of L2 English learners. We introduce ProLex, a novel benchmark designed to assess systems’ ability to generate appropriate and language proficiency-oriented substitutes for given target words in the context. Besides, we also fine-tuned open-source language models on task-specific synthetic data, achieving results that are on par with, or better than GPT-4 and ChatGPT. In future work, we will iteratively expand ProLex by leveraging a larger source corpus and incorporating a more comprehensive set of substitutes.
@@ E.2.1 Synthesizing task-specific training data
We start by selecting the context sentences from the TOEFL-11 dataset. Concretely, we randomly select $500$ sentences from each proficiency level [^11] Note that there are three proficiency levels in the TOEFL-11 dataset, namely high, medium and low., rendering a total of $1,500$ sentences. We also make sure that the selected sentences are not in ProLex. Similar to section 3.1, we perform grammar error correction to the sentences extracted from low and medium-level essays, while sentences from high-level essays are retained in their original form. Subsequently, for each sentence, we prompt GPT-4 to select a target word based on a randomly pre-defined part-of-speech tag and generate the proficiency-oriented substitutes. The complete prompt for this process is shown in Table 13. Specifically, we provide five in-context examples to guide the generation process. As a result, we take the final substitutes generated for each sentence as the proficiency-oriented substitutes. The resulting synthesized data is post-processed to fine-tune Vicuna 7B/13B and Llama-2 7B/13B models. In total, after post-processing [^12] The post-processing involves dropping null values and duplicated rows., the synthesized dataset contains $1,383$ unique $(w,s)$ pairs, with $6,375$ candidate substitutes and $4,982$ final substitutes. The part-of-speech tags for all selected target words are uniformly distributed, with $361$ nouns, $335$ verbs, $357$ adjectives, and $330$ adverbs.

### v3

@@ Abstract
Lexical Substitution discovers $appropriate$ substitutes for a given target word in a context sentence. However, the task fails to consider substitutes that are of equal or higher proficiency than the target, an aspect that could be beneficial for language learners looking to improve their writing. To bridge this gap, we propose a new task — language proficiency-oriented lexical substitution. We also introduce ProLex, a novel benchmark designed to assess systems’ ability to generate not only appropriate substitutes but also substitutes that demonstrate better language proficiency. Besides the benchmark, we propose models that can automatically perform the new task. We show that our best model, a Llama2-13B model fine-tuned with task-specific synthetic data, outperforms ChatGPT by an average of 3.2% in F-score and achieves comparable results with GPT-4 on ProLex. [^1] Data and code available: https://github.com/BillyZhang24kobe/LS_Proficiency
@@ 1 Introduction
The beginner L2 learners can leverage existing lexical substitution systems Zhou et al. (2019); Lacerra et al. (2021); Yang et al. (2022); Wada et al. (2022); Qiang et al. (2023) to enhance their vocabulary breadth. These systems are designed to identify contextually suitable substitutes for a target word, thereby assisting learners in discovering appropriate substitutions. However, merely knowing these contextually appropriate substitutes is insufficient for L2 learners. Prior benchmarks McCarthy and Navigli (2007); Kremer et al. (2014); Horn et al. (2014); Lee et al. (2021) for lexical substitution focus solely on the contextual appropriateness of generated substitutes. To enhance the vocabulary diversity and writing proficiency of English learners, guided by the principles of the zone of proximal development Vygotsky (1978), we propose that the substitutions should also reflect an equal or higher level of language proficiency compared to the original target word.
In this work, we present ProLex, a novel lexical substitution benchmark that evaluates system performances on language proficiency-oriented lexical substitution, a new task that proposes substitutes that are not only contextually suitable but also demonstrate advanced-level proficiency. To construct ProLex, we begin by selecting target words and sentences according to their frequency in TOEFL-11 Blanchard et al. (2013), a comprehensive corpus of essays written by non-native English learners. This approach ensures a data distribution that is more representative of L2 English learners. Subsequently, to effectively harness human expertise in annotating advanced proficiency-oriented lexical substitutions, we adopt the methodology recommended by Lee et al. (2021) — directly having humans judge the appropriateness of candidate substitutes instead of generating them ad hoc. To facilitate this, we employ GPT-4 to generate an initial pool of substitute candidates. Then, the human annotators assess these candidates based on their contextual appropriateness using our language-learning-oriented annotation scheme. The proposed scheme takes into account the preservation of semantic meaning, common English collocations, lexical diversity, and grammatical correctness. We specifically recruited two annotators who are pursuing their PhDs in Linguistics, specializing in English language teaching and education. Following human annotation, we use the Common European Framework of Reference (CEFR) Council of Europe. Council for Cultural Co-operation. Education Committee. Modern Languages Division (2001) to remove substitutions that have lower proficiency levels compared to the original target word.
@@ 2 Related Work
The large-scale non-native English writing corpus, TOEFL-11 Blanchard et al. (2013), contains writing samples from non-native TOEFL test takers from 2006 to 2007. In particular, it contains $12,100$ essays, each of which is categorized into one of the three score levels: high, medium, and low. Focusing on enhancing the writing skills of English learning beginners, our methodology involves selecting context sentences and target words primarily from essays with medium and low scores. This approach is based on the assumption that beginner L2 English learners are likely to possess a more limited vocabulary range Fan (2000).
@@ 3.1 Data Creation
The TOEFL-11 dataset comprises $12,100$ essays, each scored as high, medium, or low. Specifically, there are $4,202$ high-scored essays, $6,568$ with medium scores, and $1,330$ categorized as low. Given the limited vocabulary range of low and medium-proficiency L2 English learners, we intend to select the words that are frequently used in their essays as the target words in our benchmark. For each essay of interest, we select mainly the content words (i.e. nouns, verbs, adjectives, and adverbs) that are used at least three times in the essay. Then, for each such target word, we randomly choose one sentence from all sentences containing that target word. The general statistics of all target words and context sentences are presented in Appendix C Table 5. In sum, we selected $2,531$ unique word-context $(w,s)$ pairs from low and medium-level essays from TOEFL-11 corpus.
@@
We observed that the chosen sentences frequently contain fundamental grammar mistakes, such as spelling errors, which can obscure the intended meaning of the context. Example grammar errors are demonstrated in appendix D. To correct these errors, we follow Rothe et al. (2021) and fine-tune a GPT-2 model Radford et al. (2019) on CLang-8 dataset. [^2] https://github.com/ google-research-datasets/clang8 The experiment details are also shown in appendix D. The resulting model is deployed to correct the basic grammar errors in all of our $2,531$ selected context sentences $s$. We then sample $1,125$ $(w,s)$ pairs for human annotations. [^3] We used $125$ pairs for calculating inter-annotator agreement and $1,000$ for the annotation. We will include more pairs to expand the benchmark in future work.
@@ 4 Dataset annotation process and statistics
Considering the necessity for substantial semantic understanding and knowledge of English collocations in the annotation process, we recruited two annotators currently pursuing PhDs in Linguistics, with specializations in English language teaching and education. On the $125$ $(w,s)$ pairs, namely $625$ $(w,s,w^{\prime})$ triples, the two annotators reached an inter-annotator agreement of $\kappa=0.60$, indicating a near-substantial level of agreement. [^5] We perform Cohen’s Kappa Cohen (1960) to calculate the inter-annotator agreement. We exclude the labels with 0 to consider only the contexts that are conceivable to both annotators (i.e. labeled as 1 or -1). Then, they annotated the complete test set separately, with each annotating $2,500$ $(w,s,w^{\prime})$ triples sampled from section 3.1. In total, among all $5,000$ candidate substitutes $w^{\prime}$, $39\%$ is labeled as 1 (accept), while $51\%$ and $10\%$ are annotated as -1 (reject) and 0 (uncertain), respectively. This entails that GPT-4 generated candidates may not always be valid in ProLex, indicating future research opportunities in using LLMs for generating semantically correct, collocationally appropriate English lexical substitutions.
@@
We further process the annotations to compose the final quadruplets $(w,s,w^{a},w^{a}_{p})$ in ProLex, where $w^{a}$ denotes a list of acceptable substitutes for $w$, and $w^{a}_{p}$ denotes the list of acceptable substitutes after filtering based on CEFR levels. We exclude the target words that have zero acceptable substitutes. Table 1 provides an overview of the statistics for the resulting data, grouped by low and medium-level proficiency. In total, there are $680$ word-sentence $(w,s)$ pairs in ProLex with at least one acceptable substitute. [^6] The Part-Of-Speech distribution for the target word $w$ in ProLex is shown in Table 6. On average, for each $(w,s)$ pair, there are $2.9$ acceptable substitutes $w^{a}$ and $2.3$ proficiency-oriented substitutes $w^{a}_{p}$. Furthermore, we provide the distribution of CEFR levels of the target words in both low and medium sentences, and the result is shown in Figure 2. As evident, a significant proportion of the chosen target words are from A1 and A2 levels, comprising 75% in low-level sentences and 55% in medium-level sentences, respectively. Moreover, in Figure 3, we also present the CEFR level distributions for acceptable substitutes $w^{a}$ and proficiency-oriented substitutes $w^{a}_{p}$. The distribution resonates closely with our research objective of ProLex, which aims to propose acceptable substitutions of high-proficiency level words for commonly used low-proficiency level words in L2 English sentences.
@@ 5.2 Evaluation Metrics
Models BERT-LS Para-LS Llama-2 Vicuna-1.5 ChatGPT GPT-4 7B 13B 7B 13B $D_{S}$ $D_{L}$ $D_{S}$ $D_{L}$ $D_{S}$ $D_{L}$ $D_{S}$ $D_{L}$ $zero$ $32$ $zero$ $32$ Soft $F^{10}$ 25.2 30.2 44.1 48.6 44.4 51.1 43.5 48.0 46.2 52.1 50.9 51.5 50.5 54.7 $F^{10}_{p}$ 20.3 25.3 39.1 43.2 38.6 46.2 38.6 43.1 40.7 46.8 45.2 45.9 44.6 48.8 Hard $F^{10}$ 20.0 20.2 31.2 46.5 27.1 49.0 31.4 45.6 32.2 49.6 47.8 48.3 48.8 52.7 $F^{10}_{p}$ 15.8 17.2 27.8 41.2 23.9 44.1 27.7 40.9 28.6 44.5 42.5 43.0 43.3 46.8
@@
Models Vicuna-1.5 Llama-2 7B 13B 7B 13B Soft $F^{10}$ 49.9 52.4 49.7 54.2 $F^{10}_{p}$ 45.0 47.0 44.6 48.4 Hard $F^{10}$ 48.1 50.5 48.0 51.6 $F^{10}_{p}$ 43.0 45.2 42.8 45.9
@@ 5.3 Baselines
As instruction tuning has proliferated in building powerful instruction following models Ouyang et al. (2022); Bai et al. (2022); OpenAI (2023); Chiang et al. (2023), we experiment with two instruction-tuned large-scale language models, namely Vicuna [^8] We used Vicuna v1.5 provided by Zheng et al. (2023). Chiang et al. (2023) and Llama-2 Touvron et al. (2023). Note that we mainly focus on the smaller variants: Vicuna 7B/13B and Llama-2 7B/13B. Since there is no existing training data for our language proficiency-oriented lexical substitution task, we adopt two approaches to synthesize the data: 1) directly generate data with GPT-4 through prompting, and 2) modify the existing lexical substitution benchmark, Swords. The prompts we employed to synthesize the data with GPT-4 are shown in section E.2.1. As for the second approach, we considered only the acceptable substitutes (i.e. score greater than 50%) from the Swords dataset and filtered the substitutes based on their CEFR levels, by only keeping the ones that have equal or higher proficiency levels compared to the target word. As a result, we denote the training data synthesized with GPT-4 as $D_{L}$, the data modified from Swords as $D_{S}$, and the aggregate of these two datasets as $D_{LS}$.
@@
Leveraging the potent zero-shot capabilities of Large Language Models (LLMs), we evaluate their performance on ProLex by prompting two widely recognized LLMs, GPT-4 and ChatGPT, [^9] Note that we used the version gpt-3.5-turbo-1106 to generate potential substitution candidates. Moreover, we also perform in-context learning with examples selected from $D_{L}$. [^10] We randomly select $32$ examples from $D_{L}$ for in-context learning with LLMs See appendix E.1 for more details.
@@ 6 Conclusion and Future Work
We propose a new task language proficiency-oriented lexical substitution to improve the vocabulary diversity and writing proficiency of L2 English learners. We introduce ProLex, a novel benchmark designed to assess systems’ ability to generate appropriate and language proficiency-oriented substitutes for given target words in the context. Besides, we also fine-tuned open-source language models on task-specific synthetic data, achieving results that are on par with, or better than GPT-4 and ChatGPT. In future, we will expand ProLex by leveraging larger L2 English writing corpora and incorporating more comprehensive sets of substitutes.
7 Limitations
Benchmark Size and Coverage
Compared to previous Lexical Substitution benchmarks McCarthy and Navigli (2007); Kremer et al. (2014); Lee et al. (2021), the size of ProLex is relatively small. However, considering that ProLex draws its data from the TOEFL-11 corpus, it could be significantly extended by incorporating sentences from additional sources authored by L2 speakers.
In addition, the systems may produce valid substitutes that are not present in ProLex during evaluation, indicating that ProLex has limitations in coverage. By employing the annotation scheme described in Section 3.2, we can iteratively update and enhance ProLex with additional gold standard substitutes, thereby expanding its coverage in future versions.
Limits of CEFR Checker
The CEFR Checker CathovenAI (2023) we used in this work is capable of assigning CEFR levels only at the word level. However, we noticed that in some cases, phrases could also serve as appropriate substitutes. Hence, extending the CEFR Checker’s functionality to include CEFR-level assignments at both the word and phrase levels would offer a more holistic evaluation in future versions of ProLex.
Phrases and Multi-word Substitutions
In this work, we posit that the integration of multi-word expressions, such as idiomatic phrases, can enhance English language proficiency Thyab (2016); Yunus and Hmaidan (2021); Al-Khawaldeh et al. (2016). Therefore, we treat all acceptable phrasal substitutes as valid substitutes for demonstrating better language proficiency compared to the single target word. In future research, we intend to investigate the methods for assessing proficiency levels in the use of phrases and multi-word expressions.
8 Ethics Statement
Reproducibility
In this work, our data creation process utilized GPT-4. We also used ChatGPT for evaluation purposes. Although they are not open-sourced language models, to facilitate the reproducibility of our results, we demonstrated all of prompts used in our paper. In addition, all the other models used in this research, are publicly available in peer-reviewed articles and referenced in this paper. All datasets, including our synthetic fine-tuning dataset and all annotated test data, will be released.
Biases
Our models are built over Vicuna and Llama-2. In this work, we did not explicitly handle any bias that exists in the two pre-trained language models.
Human Annotators
Both annotators were recruited from the doctoral programs in the linguistics department, and they specialize in English language teaching and education. They were paid at a rate of \$13 per hour. To ensure the privacy and anonymity of all contributors, no personal or demographic information was collected.
9 Acknowledgement
We would like to thank Qingyang Wu, Kun Qian, Siyan Li, Matthew Toles, Yu Li, and Xiao Yu for their valuable discussions and suggestions on the paper. We also want to express our gratitude to Cathoven AI for providing free access to their CEFR checker APIs. In addition, we thank our expert annotators for their time and contributions to the completeness of the benchmark.
@@ E.2.1 Synthesizing task-specific training data
We start by selecting the context sentences from the TOEFL-11 dataset. Concretely, we randomly select $500$ sentences from each proficiency level, [^11] Note that there are three proficiency levels in the TOEFL-11 dataset, namely high, medium and low. rendering a total of $1,500$ sentences. We also make sure that the selected sentences are not in ProLex. Similar to section 3.1, we perform grammar error correction to the sentences extracted from low and medium-level essays, while sentences from high-level essays are retained in their original form. Subsequently, for each sentence, we prompt GPT-4 to select a target word based on a randomly pre-defined part-of-speech tag and generate the proficiency-oriented substitutes. The complete prompt for this process is shown in Table 13. Specifically, we provide five in-context examples to guide the generation process. As a result, we take the final substitutes generated for each sentence as the proficiency-oriented substitutes. The resulting synthesized data is post-processed to fine-tune Vicuna 7B/13B and Llama-2 7B/13B models. In total, after post-processing, [^12] The post-processing involves dropping null values and duplicated rows. the synthesized dataset contains $1,383$ unique $(w,s)$ pairs, with $6,375$ candidate substitutes and $4,982$ final substitutes. The part-of-speech tags for all selected target words are uniformly distributed, with $361$ nouns, $335$ verbs, $357$ adjectives, and $330$ adverbs.